# SplatOverflow: Asynchronous Hardware Troubleshooting

### v2

@@ 4. Related Work
SplatOverflow builds on studies of hardware maintenance practices, hardware documentation support, telepresence for expert support, and spatial rendering techniques.
@@ 4.1. Hardware Maintenance Practice
While studying the practices of Xerox repair technicians, Julian Orr details how communication between expert technicians, end-users, and the machine plays an essential role in making sense of hardware issues (ORR, 1996a). Orr underscores how hardware maintenance cannot exhaustively be preempted in the hardware’s documentation, instead describing the improvisational nature of diagnosing hardware issues. Rather than focusing on design and innovation when considering technological progress, Jackson (2014) introduces ”broken world thinking” and argues for foregrounding repair and maintenance practices. Similar ethnographic accounts of repair and “fixer” communities have examined commonalities and differences in these groups’ norms, goals, and processes (Houston et al., 2016; Jackson and Kang, 2014). More recently, Subbaraman and Peek (2023b) argue that maintenance is a core component of using digital fabrication workflows in 3D printing communities. Like Orr, they caution against addressing maintenance needs with fully automated procedures and instead advocate for systems to be designed with maintenance in mind.
4.2. Hardware Documentation Support
Documentation for assembly instructions, usage guides and maintenance tasks play an integral role in how users interact with and navigate hardware. Such documentation is distinct from the files that define the hardware, but researchers have argued that they contribute just as significantly to the hardware (Bonvoisin et al., 2017; ORR, 1996b). In HCI, researchers have examined how such documentation may be made easier to produce. Mariscal-Melgar et al. (2023) propose a semi-automated method for generating and updating assembly documentation from CAD software. Milara et al. (2019) propose software tools for makers to create documentation as they are designing. Agrawala et al. (2003) describe techniques for designing assembly instructions that are easy for end-users to understand. Similarly, researchers have explored how to make it easier for end-users to explore documentation. MagicNeRFLens(Li et al., 2023) allows users to navigate a NeRF scene in virtual reality, with additional documentation overlaid onto the virtual scene. ARDW (Chatterjee et al., 2022) is an augmented reality projection workbench that overlays documentation onto a PCB to support debugging.
@@ 4.3. Telepresence for Expert Support
In software development (Anderson et al., 2012) and document editing (Posner et al., 1993), all collaborators typically access the same source code instead of compiled or high-level artifacts. Direct access to the collaborative artifact enables asynchronous support across the web (Anderson et al., 2012). HCI researchers examining synchronous support have proposed multiple theories and developed speculative systems for how remote experts can virtually occupy a shared task space with non-expert users to provide guidance (Buxton, 1992; Oda et al., 2015; Gauglitz et al., 2012; Fussell et al., 2000a). Telepresence solutions often try to establish a shared reference space, in which participants can communicate with one another with deictic expressions and gestures as they would if they were in person (Johnson et al., 2021; Gutwin and Greenberg, 2002; Dourish and Bellotti, 1992). Such a reference space depends on both participants being able to communicate within a common context. In a survey on augmented reality systems for collocated experiences, Radu et al. (2021) found that collaborators need a shared environment to ground their communication. Moreover, they find that collaborators need tools to guide attention and give instructions by referencing objects in the shared task space. Chastine et al. (2006) develop a formal model of such inter-referential awareness, providing a detailed framework for how designers can develop referential systems.
@@
Advances in augmented and virtual reality have led to work studying how experts can remotely collaborate on physical tasks (Wang et al., 2021). Researchers have proposed systems for ’mentoring’ tasks, where one less experienced user seeks out guidance from an expert user who views the scene remotely using mobile AR to bring the physical environment of users to remote collaboration (Fussell et al., 2000b; Johnson et al., 2021; Gauglitz et al., 2012). Systems may have experts view the perspective of remote users in VR and communicate guidance through overlaid annotations (Thoravi Kumaravel et al., 2020; Chen et al., 2021; Yu et al., 2022).
@@
Alternatively, some systems capture 3D reconstructions (Izadi et al., 2011) and allow remote collaborators to interact by combining with teleoperated robots as in VRoxy (Sakashita et al., 2023) or by integrating live 360°video as Teo et al. (2019) demonstrated.
@@
In these systems, the interaction is synchronous, and the guidance provided is ephemeral, making it difficult for third-parties to replay and observe at a later time. Heimdall (Karchemsky et al., 2019) is an alternative approach to remote collaboration for prototyping electrical circuits featuring a bespoke scanning system. Instructors can navigate around the prototype circuit and inspect its schematics through an instrumented breadboard. This enables instructors to remotely examine and debug student circuits without relying on a student to navigate around their breadboard.
4.4. Spatial Rendering
Neural Radiance Fields (NeRFs) introduced a novel representation for 3D scenes that captured high-quality geometry estimations and view-dependent textures from collections of static images (Mildenhall et al., 2020). This work builds on earlier research in capturing real-world objects using images from multiple viewpoints (McMillan and Bishop, 1995; Chen and Williams, 1993; Penner and Zhang, 2017; Wiles et al., 2020) and scenes from images captured in-the-wild (Snavely et al., 2006). Müller et al. (2022) speed up training and rendering times for NeRFs and enable rapid capture and real-time navigation through the 3D scenes. More recently, researchers used collections of 3D Gaussian distributions to describe scenes in a process called 3D Gaussian Splatting (Kerbl et al., 2023). A splat represents the scene statically and can be rendered in real-time using traditional graphics pipelines. These approaches build on COLMAP, a structure-from-motion (SfM) tool that constructs a point cloud and estimates camera pose from image frames (Schönberger and Frahm, 2016; Schönberger et al., 2016). SplatOverflow is generally agnostic to scanning technology but leverages 3D Gaussian Splatting to capture the end-user’s workspace due to its fast training time and real-time rendering capabilities.
@@ 5. SplatOverflow
This section describes the design of SplatOverflow, shown in figure 10. We first discuss the constituent artifacts that comprise a SplatOverflow scene, including the benefits and shortcomings of each artifact. Then, we explain the construction of the SplatOverflow scenes, discuss how it facilitates asynchronous communication using SplatOverflow gestures, and finally describe how data from issues can be indexed and retrieved using a CAD model.
@@ 5.1. Constituent Artifacts
SplatOverflow draws upon the benefits of a few existing artifacts that can be used to describe hardware. At a minimum, SplatOverflow relies on a scan of the hardware and a CAD model that describes the design of the hardware. Both artifacts have benefits and drawbacks as boundary objects for communicating the details and instructions to troubleshoot hardware. SplatOverflow combines these artifacts to construct a single boundary object that amplifies the benefits and minimizes the drawbacks of each constituent artifact.
@@ 5.1.1. CAD Models
CAD models can precisely and accurately describe the design of hardware. They also offer various viewing, selection, and manipulation tools to help make sense of hardware issues. For example, CAD models can let designers inspect internal mechanisms that aren’t visible in the assembled hardware. This ability to peer through parts can help designers reason about mechanical issues without taking apart the hardware. These models can also be the basis of generating visual assembly instructions that are easy for users to follow (Agrawala et al., 2003; Mariscal-Melgar et al., 2023).
@@ 5.1.2. Scans
The drawbacks of relying on scans are that they only capture visible geometry and do not know the semantic meaning of what they are capturing. Most scanning methods cannot safely penetrate materials to scan internal components. As a result, scans capture only an outer shell of the hardware geometry. Moreover, as scanning methods often do not know what they are scanning, it is challenging to segment out different parts. These two drawbacks are related in that they highlight how references are hard to anchor to scan data. An asynchronous workflow using scans must overcome this referential ambiguity.
@@ 5.2. A SplatOverflow Scene
SplatOverflow scenes comprise two essential artifacts: a scan of the user’s workspace aligned and registered onto a CAD model of the hardware. Our implementation generates the scans using 3D Gaussian Splatting (Kerbl et al., 2023). We chose Gaussian Splatting as it can generate high-quality scans rapidly using only user-captured video data as input. We use open-source CAD models saved as .glb files with additional data to capture the assembly constraints in the model. The 3D Gaussian Splat (referred to here as splat) is aligned and registered to the CAD model using aruco markers that are placed on the hardware at known locations (Munoz-Salinas, 2012; Olson, 2011). The aligned artifacts support selecting and manipulating individual components on the hardware segmented using the CAD model and inspecting the physical state of the hardware as captured in the splat. SplatOverflow scenes are implemented in WebGL and run in the browser. As a result, they can be shared with anyone online, requiring no installation or OS-specific set-up. Scenes can be viewed on personal computers, mobile devices, and AR/VR headsets that support WebXR.
@@ 5.2.1. Scanning a Workspace
Scanning the hardware and its workspace captures the as-built state of the hardware. This includes modifications, job configurations, or parts not traditionally included in CAD, such as wires or cables. It can also capture the context in which the hardware is used. Scanning the physical hardware should ideally be fast and produce a high-fidelity rendering without requiring specialty hardware. To this end, SplatOverflow uses 3D Gaussian Splatting (Kerbl et al., 2023) to scan the hardware and workspace.
The input to the scan is a video of the user’s workspace. This video should adequately capture the hardware from multiple angles and include footage of any areas the user wants to troubleshoot. We implement a mobile phone interface for SplatOverflow to guide the user in creating a high-quality splat. Figure 2 shows this mobile interface in use.
@@
The video is then sampled in two passes. First, at a fixed frame rate (4 FPS in our examples) and then again to more densely sample frames containing aruco tags. These two passes generate a set of images, I. SplatOverflow then feeds this set of images to COLMAP (Schönberger and Frahm, 2016; Schönberger et al., 2016) to generate an intermediate Structure-from-Motion model of the local users’ workspace. This model, M, includes a coarse point cloud representation of the workspace and camera pose estimates for each registered frame from set I. SplatOverflow uses model M to generate a splat of the recorded scene.
SplatOverflow takes roughly four minutes to generate a 3D Gaussian Splat from an input video filmed in 1080p resolution. This allows the local user to access SplatOverflow’s features quickly. SplatOverflow generates a lower-quality 3D Gaussian Splat from a downsampled version of the input video (at 720p resolution) to further accelerate processing. This splat is slightly worse at capturing details but still lets the local user query documentation and inspect past issues. The lower-quality splat is swapped out for a higher-quality one when processing is completed on the server. This low-resolution preview speeds up the feedback loop when creating a SplatOverflow scene.
@@ 5.2.2. Localizing Video Feeds
While 3D Gaussian Splats can capture the static details of a workspace with high resolution, they do not capture dynamic details in a scene. This is important for communicating the nature of many hardware issues, such as play or slop in an assembly. SplatOverflow allows users to capture these dynamic details as short video clips and aligns these clips in the SplatOverflow scene using an estimate of the camera pose at each frame. This allows remote maintainers to review the perspective from which a video was shot as well as the content of the video. Figure 5 shows the video feed in the SplatOverflow scene.
SplatOverflow estimates the pose of a video feed using COLMAP’s utility to localize frames within an existing model (Schönberger and Frahm, 2016; Schönberger et al., 2016), in our case M. Given an input video, V, SplatOverflow samples frames at a fixed frame rate (e.g., 2 FPS) and uses COLMAP to localize each frame within the SfM scene (Schönberger and Frahm, 2016; Schönberger et al., 2016). The poses returned from COLMAP are used to place the video into the SplatOverflow scene.
5.2.3. Registering a Gaussian Splat to a 3D CAD Model
Our contribution is unlocked by creating a novel boundary object, a SplatOverflow scene, that can capture the physical state of the local user’s hardware and anchor instructions made by remote maintainers. At the core of a SplatOverflow scene is a 3D Gaussian Splat registered onto a CAD model of the hardware. Fusing these two artifacts allows references to be linked to common digital artifacts and render instructions in the local user’s environment.
SplatOverflow aligns and registers a splat of the local user’s hardware to the 3D CAD model in two steps using aruco tags placed in known locations on the hardware (Munoz-Salinas, 2012). SplatOverflow first uses the SfM model M to compute a three-dimensional coordinate of every corner of each aruco tag (Meyer, 2023). These corners are used to re-scale the splat to be the same size as the CAD model. Next, SplatOverflow uses Arun’s method (Arun et al., 1987) to compute a rigid transformation going from corners in the SfM model’s coordinate space to CAD coordinate space. If there were no moving parts in the CAD assembly, this transformation yields an aligned and registered splat. However, hardware will often have moving parts. These degrees of freedom mean that the hardware can be in one of many states, which are unlikely to match the static state of the CAD model. To address this, SplatOverflow recommends at least one constaint tag be placed on each degree of freedom, and one grounding tag be placed on a rigid part of the hardware. When a CAD model is being prepared, the model’s maintainer indicates which tag ID maps to which degree of freedom. With this precondition, SplatOverflow uses the same detection method to find the centers of each constaint tag and computes an offset from the grounding tag to the constraint tag. This offset is then used to calculate a transformation that automatically aligns the CAD component to the position of the part in real life. Following these two steps, the hardware parts are aligned to their corresponding parts in the CAD model.
5.2.4. Post Processing a Splat Using Features of the CAD Model
Once the splat has been aligned and registered, SplatOverflow leverages information from the CAD model to post-process the splat. Firstly, SplatOverflow can separate captured points on the hardware from background points. SplatOverflow uses a signed distance function from points in the splat to CAD meshes. SplatOverflow not only allows users to toggle the subject and background on and off, but it enables the post-processing of a splat to preserve the local user’s privacy. By default, SplatOverflow removes the background details from a hardware scene and shares only parts of the splat that correspond to the hardware or the workspace the hardware rests on. Figure 11 shows a splat as it is generated and the same splat after SplatOverflow’s privacy filter. This has the benefit of protecting the privacy of local users while also cleaning up floating artifacts that can be generated when training a splat.

### v3

@@ 4. Related Work
SplatOverflow builds on studies of hardware maintenance practices, documentation systems, telepresence systems for expert support, and spatial rendering techniques.
@@ 4.1. Hardware Maintenance Practice
SplatOverflow draws on the findings of Julian Orr’s ethnographic work, which studies the sharing practices of Xerox repair technicians in the field (ORR, 1996). Orr’s work found that such troubleshooting was improvisational and centered collaborative sensemaking between a technician, a client, and a machine (ORR, 1996). This form of collaboration between different parties remains essential for effective maintenance in modern, distributed communities of hardware users. Dunn et al. (2023) study one such distributed community: users of the Jubilee motion platform, an open-source hardware project. They find that encouraging community-driven maintenance enabled the hardware to evolve to address design issues that would likely have gone unaddressed otherwise (Dunn et al., 2023). Orr emphasizes how hardware maintenance and troubleshooting practice cannot exhaustively be preempted through documentation (ORR, 1996). The practical nature of hardware issues is that they are non-canonical and can only be addressed through pools of shared knowledge that continue to evolve over time (Bobrow and Whalen, 2002; Brand, 2025).
@@
More recently, Subbaraman and Peek (2023b) argue that maintenance is a core component of using digital fabrication workflows in 3D printing communities. Like Orr, they caution against addressing maintenance needs with fully automated procedures and instead advocate for systems to be designed with maintenance in mind. This resembles Jackson’s (Jackson, 2014) concept of ”broken world thinking,” which argues for foregrounding repair and maintenance practices when considering technological progress. SplatOverflow revisits these ideas to examine how such troubleshooting could be coordinated asynchronously and how systems supporting hardware maintenance can contribute workflows to pools of practical knowledge in hardware communities.
4.2. Documentation Systems
Documentation for assembly instructions, usage guides, and maintenance tasks is integral to how users interact with and navigate hardware. Such documentation is distinct from the design files that define the hardware, but researchers have argued that they contribute just as significantly to its success and adoption (Bonvoisin et al., 2017; ORR, 1996). To this end, HCI researchers have examined how hardware designers can efficiently create documentation that is useful to end-users. Mariscal-Melgar et al. (2023) propose a semi-automated method for generating and updating assembly documentation by leveraging data from the hardware’s CAD model. This lets designers compile new assembly instructions as designs change and keep documentation in sync with updates to the hardware. Milara et al. (2019) propose software tools for makers to create documentation as they work on projects to ensure that essential design decisions, lessons from failures, and the process narrative are not lost. Tran O’Leary et al. (2024) demonstrate how to create fabrication workflows in a manner that foregrounds reproducibility by others. By interweaving digitally controlled processes and physical interventions when defining a fabrication workflow, Tandem guides users through the manual steps critical to successfully reproducing a workflow (Tran O’Leary et al., 2024).
Researchers have also explored how to make it easier for end-users to explore existing technical documentation. For instance, MagicNeRFLens (Li et al., 2023) allows users to navigate a NeRF scene in virtual reality, with additional documentation overlaid onto the virtual scene. ARDW (Chatterjee et al., 2022) is an augmented reality projection workbench that overlays documentation onto a PCB to support debugging. These projects present overlay and interaction techniques that address challenges associated with context shifting between design, documentation, and physical instantiation of hardware. This research underscores that the successful documentation, distribution, and reproduction of hardware and hardware workflows extends far beyond sharing digital design files. However, these projects demonstrate that there is significant value in incorporating connections between digital design files, technical documentation, and physical instantiations of hardware. SplatOverflow builds on this work by leveraging existing CAD model information to author, share, visualize, and index hardware troubleshooting workflows.
@@ 4.3. Telepresence for Expert Support
In software development (Anderson et al., 2012) and document editing (Posner et al., 1993), all collaborators typically access the same source code instead of compiled or high-level artifacts. Direct access to the collaborative artifact enables asynchronous support across the web (Anderson et al., 2012).
@@
HCI researchers examining synchronous support have proposed multiple theories and developed speculative systems for how remote experts can virtually occupy a shared task space with non-expert users to provide guidance (Buxton, 1992; Oda et al., 2015; Gauglitz et al., 2012; Fussell et al., 2000a). Telepresence solutions often try to establish a shared reference space, in which participants can communicate with one another with deictic expressions and gestures as they would if they were in person (Johnson et al., 2021; Gutwin and Greenberg, 2002; Dourish and Bellotti, 1992). Such a reference space depends on both participants being able to communicate within a common context. In a survey on augmented reality systems for collocated experiences, Radu et al. (2021) found that collaborators need a shared environment to ground their communication. Moreover, they find that collaborators need tools to guide attention and give instructions by referencing objects in the shared task space. Chastine et al. (2006) develop a formal model of such inter-referential awareness, providing a detailed framework for how designers can develop referential systems.
@@
Advances in augmented and virtual reality have led to work studying how experts can remotely collaborate on physical tasks (Wang et al., 2021). Researchers have proposed systems for ’mentoring’ tasks, where one less experienced user seeks out guidance from an expert user who views the scene remotely using mobile AR to bring the physical environment of users to remote collaboration (Fussell et al., 2000b; Johnson et al., 2021; Gauglitz et al., 2012). Some systems let experts view the perspective of remote users in VR and communicate guidance through overlaid annotations (Thoravi Kumaravel et al., 2020; Chen et al., 2021; Yu et al., 2022).
@@
Alternatively, systems also capture 3D reconstructions (Izadi et al., 2011) and allow remote collaborators to interact by combining with teleoperated robots as in VRoxy (Sakashita et al., 2023) or by integrating live 360°video as Teo et al. (2019) demonstrated.
@@
In these systems, the interaction is synchronous, and the guidance provided is ephemeral, making it difficult for third parties to replay and observe at a later time. Heimdall (Karchemsky et al., 2019) is an alternative approach to remote collaboration for prototyping electrical circuits featuring a bespoke scanning system. Instructors navigate around the prototype circuit and inspect its schematics through an instrumented breadboard. This enables instructors to remotely examine and debug student circuits without relying on a student to navigate around their breadboard. SplatOverflow marries insights from synchronous remote expert support systems with the flexibility and observability enabled by asynchronous troubleshooting.
4.4. Scanning and Spatial Rendering
To troubleshoot hardware issues, SplatOverflow must be able to capture a local user’s hardware, ideally without specialized equipment. To do this, it leverages research in scanning and scene representation techniques. Recently, there has been substantial progress in novel-view synthesis techniques that can construct a 3D scene from collections of posed 2D images (Mildenhall et al., 2020; Müller et al., 2022; Kerbl et al., 2023). These approaches build on COLMAP, a structure-from-motion (SfM) tool that constructs a coarse point cloud and estimates camera pose from image frames (Snavely et al., 2006; Schönberger and Frahm, 2016; Schönberger et al., 2016). These tools and techniques allow SplatOverflow users to capture high-fidelity scans using commodity smartphone cameras they likely already own.
Although SplatOverflow is generally agnostic to scanning technology, our implementation leverages 3D Gaussian Splatting (Kerbl et al., 2023) to capture a workspace due to its fast training time and real-time rendering capabilities. A splat represents the scene as a collection of 3D Gaussian distributions and can be rendered in real-time using traditional graphics pipelines. Notably, there is considerable progress in reducing training times and improving the visual definition of Gaussian splats (Guédon and Lepetit, 2023; Liu et al., 2024; Wang et al., [n. d.]). These methods enable fast 3D scene reconstruction with accurate geometric information from limited data, lowering the overhead of capturing a scene for users.
@@ 5. SplatOverflow
This section describes the design of SplatOverflow. We discuss the constituent artifacts that comprise a SplatOverflow scene, including the benefits and shortcomings of each artifact. Then, we explain the construction of the SplatOverflow scenes, how it facilitates asynchronous communication using gestures, and how data from issues can be indexed and retrieved using a CAD model.
@@ 5.1. Constituent Artifacts
At a minimum, SplatOverflow requires a scan of the hardware and a CAD model that describes the hardware’s design. SplatOverflow combines these artifacts to construct a single boundary object that amplifies the benefits and minimizes the drawbacks of each constituent artifact, as well as addressing the limitations of traditional asynchronous troubleshooting artifacts described in Section 3.
@@ 5.1.1. CAD Models
CAD models can precisely and accurately describe the design of hardware. They also offer various viewing, selection, and manipulation tools to help make sense of hardware issues. For example, CAD models allow designers to inspect internal mechanisms that are not visible in the assembled hardware. This ability to peer through parts can help designers reason about mechanical issues without taking apart the hardware. These models can also be the basis of generating visual assembly instructions that are easy for users to follow (Agrawala et al., 2003; Mariscal-Melgar et al., 2023).
@@ 5.1.2. Scans
The drawbacks of relying on scans are that they only capture visible geometry and do not know the semantic meaning of what they are capturing. Most scanning methods cannot safely penetrate materials to scan internal components, so scans capture only the outer shell of the hardware geometry. Moreover, as scanning methods often do not know what they are scanning, isolating different parts in an assembly is challenging. These drawbacks are related in that they highlight how references are hard to anchor to scan data.
@@ 5.2. A SplatOverflow Scene
SplatOverflow scenes comprise two essential artifacts: a scan of the user’s workspace aligned and registered onto a CAD model of the hardware. Our implementation generates the scans using 3D Gaussian Splatting (Kerbl et al., 2023). We chose Gaussian Splatting as it can generate high-quality scans rapidly using only user-captured video data as input. We use open-source CAD models saved as .glb files with additional data to capture the assembly constraints in the model. The 3D Gaussian Splat (referred to here as splat) is aligned and registered to the CAD model using ArUco markers that are placed on the hardware at known locations (Munoz-Salinas, 2012; Olson, 2011).
The aligned artifacts support selecting and manipulating individual components on the hardware segmented using the CAD model and inspecting the physical state of the hardware as captured in the splat. SplatOverflow scenes are implemented in WebGL and run in the browser. As a result, they can be shared with anyone online, requiring no installation or OS-specific set-up. Scenes can be viewed on personal computers, mobile devices, and virtual reality headsets that support WebXR.
@@ 5.2.1. Scanning a Workspace
Scanning the hardware and surrounding workspace captures the as-built state of the hardware. This includes modifications, job configurations, or parts not traditionally included in CAD, such as wires or cables.
@@
The input to the scan is a video of the user’s workspace captured through SplatOverflow’s mobile interface, as described in Section 2. The video is then sampled in two passes. First, at a fixed frame rate (4 FPS in our examples) and then again to more densely sample frames containing ArUco tags to generate a set of images. SplatOverflow feeds this set of images to COLMAP (Schönberger and Frahm, 2016; Schönberger et al., 2016) to generate an intermediate Structure-from-Motion model of the local users’ workspace. SplatOverflow uses this model to generate a splat of the recorded scene.
SplatOverflow takes roughly four minutes to generate a 3D Gaussian Splat from a video filmed in 1080p resolution and roughly two minutes from a 360p video. Generating the splat is the primary bottleneck for developing a SplatOverflow scene. Processing times could be reduced by using more efficient splat generation techniques from recent research, such as MVSGaussian (Liu et al., 2024).
5.2.2. Registering a Gaussian Splat to a 3D CAD Model
SplatOverflow automatically aligns and registers a splat of the local user’s hardware to the 3D CAD model in two steps using ArUco tags placed in known locations on the hardware (Munoz-Salinas, 2012). SplatOverflow first uses the SfM model generated by COLMAP (Schönberger et al., 2016; Schönberger and Frahm, 2016) to compute a three-dimensional coordinate of every corner of each ArUco tag (Meyer, 2023). These corners are used to rescale the splat to be the same size as the CAD model. Next, SplatOverflow uses Arun’s method (Arun et al., 1987) to compute a rigid transformation going from corners in the SfM model’s coordinate space to CAD coordinate space. If there are no moving parts in the CAD assembly, this transformation yields an aligned and registered splat.
However, hardware will often have moving parts. These degrees of freedom mean that the hardware can be in one of many states, which are unlikely to match the static state of the CAD model. To address this, SplatOverflow recommends placing at least one constraint tag on each degree of freedom and one grounding tag on a stationary part of the hardware. When a CAD model is being prepared, the model’s maintainer indicates which tag ID maps to which degree of freedom. With this precondition, SplatOverflow uses the same detection method to find the centers of each constraint tag and computes an offset from the grounding tag to the constraint tag. This offset is then used to calculate a transformation that automatically aligns the CAD component to the position of the part in real life. Following these two steps, the hardware parts are aligned to their corresponding parts in the CAD model.
5.2.3. Post Processing a Splat Using Features of the CAD Model
Once the splat has been aligned and registered, SplatOverflow leverages information from the CAD model to post-process the splat. SplatOverflow uses a signed distance function from points in the splat to CAD meshes to identify what region of the splat contains relevant information about the hardware. By default, SplatOverflow uses this to remove the background details from a scene to preserve the local user’s privacy and share only parts of the splat that correspond to the hardware or the workspace the hardware rests on. Figure 9 shows a splat as it is generated and the same splat after SplatOverflow’s privacy filter.
Figure 9. This figure has two subfigures in it: A and B. The left subfigure A shows an unaltered Gaussian splat of a 3D printer used by SplatOverflow. The right subfigure B shows a pruned Gaussian splat of the 3D printer that is used to preserve the user’s privacy and not share background details with a community of users.
@@ 5.2.4. Localizing Video Feeds
SplatOverflow allows users to capture dynamic details of their issue as short video clips and aligns these clips in the SplatOverflow scene using an estimate of the camera pose at each frame. This allows remote maintainers to review the perspective from which a video was shot as well as the content of the video. Figure 4 shows the video feed in the SplatOverflow scene. SplatOverflow estimates the pose of a video feed by localizing individual frames into an existing COLMAP model (Schönberger and Frahm, 2016; Schönberger et al., 2016).